# Ion Temperature Anisotropy Limits from Magnetic Curvature Scattering in Magnetotail Reconnection Jets

### v1

@@ II.2.2 Speiser dominated regime
In the second regime, we consider that cold ions provide a negligible current ($n_{c}\approx 0$, $j_{cy}\approx 0$), and Speiser ions carry the current. Hence, Eq. 1 becomes
$\dfrac{n_{s}}{n_{i}}\left(A_{s}-1\right)=\left(\dfrac{T_{i\parallel}}{T_{i\bot}}-1\right)\leq 0.$ (9)
Substituting Eq. 9 into Eq. 3, the current density becomes
$j_{y}=j_{0}\left[\eta_{e}+2\left(1-\dfrac{T_{i\parallel}}{T_{i\bot}}\right)\eta_{s}\right],$ (10)
where we have used the approximation of the Speiser ion VDF [burkhart_ion_1992, burkhart_particle_1992, artemyev_ion_2019], with $A_{s}\approx 1/2$. The corresponding adiabaticity parameter becomes
$\kappa^{2}=\dfrac{1}{\beta_{i\parallel}}\dfrac{T_{i\parallel}}{T_{i\bot}}\dfrac{1}{\eta_{e}+2\left(1-\dfrac{T_{i\parallel}}{T_{i\bot}}\right)\eta_{s}}.$ (11)
To avoid the scattering of hot ions into a quasi-adiabatic Speiser regime (when the entire ion population would move with the drift speed exceeding thermal speed and the current sheet would be strongly unstable), the magnetic field curvature parameter $\kappa$ must remain close to unity [buchner_regular_1989, chen_nonlinear_1992]. Thus, $\kappa=1$ represents the marginal stability condition for the perpendicular anisotropy. Therefore, from Eq. 11, we obtain the second threshold
$\mathcal{R}_{s}(\eta_{e},\eta_{s},\beta_{i\parallel})=\dfrac{1}{1+\eta_{e}/2\eta_{s}}\left(1+\dfrac{1}{2\eta_{s}}\dfrac{1}{\beta_{i\parallel}}\right).$ (12)
In the limit $\eta_{s}\ll 1$ ($v_{sD}\ll v_{iT\bot}$), the Speiser population is embedded within the hot ion population, and the threshold reduces to $\mathcal{R}_{s}(\eta_{e},\eta_{s}\rightarrow 0,\beta_{i\parallel})\rightarrow 1/\eta_{e}\beta_{i\parallel}$. In this regime, magnetic field stretching is governed by the electron current such that for a given $\beta_{i\parallel}$ and $T_{i\bot}/T_{i\parallel}$, ions with $v_{iT\bot}>V_{Ai}/\sqrt{\eta_{e}}$ will scatter into a Speiser-type orbit. In contrast, for $\eta_{s}\gg 1$ ($v_{sD}\gg v_{iT\bot}$), the threshold approaches $\mathcal{R}_{s}(\eta_{e},\eta_{s}\rightarrow\infty,\beta_{i\parallel})\rightarrow 1$, corresponding to the most unstable configuration with a strong drift current carried by quasi-adiabatic Speiser orbits, characteristic of a drift-type instability [lapenta_kinetic_1997, daughton_unstable_1999].
@@ III Spacecraft Observations
We validate our model using in situ spacecraft observations in the Earth’s magnetotail. We employ measurements in the near-Earth magnetotail ($X_{GSM}>-29~R_{E}$, in Geocentric Solar Magnetospheric (GSM) coordinates) from the Magnetospheric Multiscale (MMS) mission [burch_magnetospheric_2016] and in the mid-tail at lunar distances ($X_{GSM}\sim-60~R_{E}$) from the Acceleration, Reconnection, Turbulence, and Electrodynamics of the Moon’s Interaction with the Sun (ARTEMIS) mission [angelopoulos_artemis_2011].
@@ III.1 Near-Earth reconnection
The MMS dataset includes 516 fast plasma flows measured between 2017 and 2021, at $X_{GSM}\geq-29~R_{E}$, in the central plasma sheet corresponding to $\beta_{i}\geq 0.5$, where $\beta_{i}=2\mu_{0}k_{B}T_{i}/B^{2}$ is the ion plasma beta [richard_are_2022]. Magnetic field measurements are obtained from the fluxgate magnetometer [russell_magnetospheric_2016], while ion and electron VDFs and their moments are provided by the Fast Plasma Investigation instrument [pollock_fast_2016]. We correct the ion VDFs for penetrating radiation using the empirical background model described in Ref. gershman_systematic_2019. To improve counting statistics and reduce uncertainties in moment calculations [gershman_calculation_2015], we average three consecutive ion VDFs over a 450 ms window. It yields a total of 235,635 ion VDFs for analysis.
@@
To illustrate the limiting cases of cold ion-dominated regime, corresponding to parallel anisotropy, and Speiser ion-dominated, corresponding to perpendicular anisotropy, we examine a reconnection outflow observed by MMS on 21 August 2017[richard_fast_2023] [Fig. 2]. During this interval, the magnetic field in Fig. 2(a) displays large-amplitude oscillations attributed to a flapping motion of the magnetotail current sheet [richard_observations_2021], so that MMS samples across the reconnection outflow shown in Fig. 2(b). Here, we rotated the magnetic field and the ion bulk velocity to the $\bm{\hat{l}}\bm{\hat{m}}\bm{\hat{n}}$ coordinates system, where $\bm{\hat{l}}=\left[0.97,\,-0.25\,0.02\right]\,\textrm{GSM}$ is the maximum variance direction obtained from maximum variance analysis (MVA) of the magnetic field with $\lambda_{max}/\lambda_{int}=15$, $\bm{\hat{n}}=(\bm{\hat{l}}\times\bm{j})/|(\bm{\hat{l}}\times\bm{j})|$, and $\bm{\hat{m}}=\bm{\hat{n}}\times\bm{\hat{l}}$ [sergeev_survey_2006]. In this coordinate system, at $|B_{l}|\approx 0$, MMS is around the center of the current sheet, while maxima in $|B_{l}|$ correspond to the current sheet edges. We present the ion temperature anisotropy $T_{i\bot}/T_{i\parallel}$ [Fig. 2(d)] with respect to the local magnetic field, where we calculated $T_{i\bot}$ and $T_{i\parallel}$ rotating the temperature tensor $\mathbf{T}_{i}$ in the magnetic field-aligned (MFA) frame [swisdak_quantifying_2016]. In this frame, one axis aligns with the local magnetic field direction $\hat{\bm{b}}$. The perpendicular directions are set to equalize the remaining diagonal components of the ion temperature tensor $\mathbf{T}_{i}=T_{i\parallel}\hat{\bm{b}}\hat{\bm{b}}+T_{i\bot}(\mathbf{1}-\hat{\bm{b}}\hat{\bm{b}})+\bm{\Theta}_{i}$, where $T_{i\parallel}$ and $T_{i\bot}$ are the parallel and perpendicular temperatures, and $\bm{\Theta}_{i}$ is the agyrotropic part with zero diagonal. We compute the magnetic field curvature radius $R_{c}$, and the corresponding adiabaticity parameter $\kappa$, using the multi-spacecraft curlometer technique [chanteur_spatial_1998, chanteur_spatial_1998a]. We find that the temperature anisotropy [Fig. 2(d)] and the adiabaticity parameter [Fig. 2(e)] vary strongly across the reconnection outflow and are inversely correlated, indicating interplay between magnetic field curvature and the ion VDFs.
@@
To examine the spatial variation of the temperature anisotropy across the current sheet, we construct profiles of $T_{i\parallel}$ and $T_{i\bot}$ as functions of the reconnecting magnetic field component $B_{l}$. We normalize $B_{l}$ by the background (lobe) magnetic field strength $B_{0}=B\sqrt{1+\beta_{i}}$, estimated from pressure balance of the magnetotail plasmasheet [asano_evolution_2003]. To facilitate comparison with numerical simulation (see Section IV), we normalize the ion temperatures $T_{i}$, $T_{i\parallel}$, $T_{i\bot}$ to the electromagnetic energy available per particle $m_{i}V_{A0}^{2}$, where $V_{A0}=B_{0}/\sqrt{\mu_{0}m_{i}n_{0}}$ is the Alfvén speed and $n_{0}$ the ion density within the current sheet. For the flapping event presented in Fig. 2, we obtain $B_{0}=21.6\,\mathrm{nT}$ and $n_{0}=0.27\,\mathrm{cm}^{-3}$, the latter computed as the average number density over intervals with $|B_{l}|/B_{0}<0.1$ well within the current sheet.
@@
Figure 3: Comparison between MMS observations and PIC simulation results for ion temperature anisotropy. Panels (a)–(d) and (e)–(h) show 1D profiles of: (a, e) ion temperature parallel to the local magnetic field, (b, f) ion temperature perpendicular to the magnetic field, (c, g) ion temperature anisotropy, and (d, h) the magnetic field curvature parameter. Black crosses represent MMS measurements during the event shown in Fig. 2. Solid blue and red lines correspond to profiles extracted along magnetic field lines in the PIC simulation, intersecting the $z=0$ plane at $\delta x=5d_{i}$ and $\delta x=11d_{i}$, respectively. Panels (i) and (j) present 2D maps of ion temperature anisotropy and curvature parameter from the PIC simulation.
@@
At the edges of the current sheet – corresponding to where $B_{l}/B_{0}=\pm 0.3$ – the ion VDFs are elongated parallel to the magnetic field, with $T_{i\bot}/T_{i\parallel}<1$, characteristic of cold, field-aligned counter streaming beams [hoshino_ion_1998, hietala_ion_2015] as described in our model. We find that instances of parallel anisotropy coincide with $\kappa>1$, indicating that the cold ions follow regular adiabatic orbits. These observations suggest that outside the current sheet center, the ion VDF primarily consists of an adiabatic, parallel-anisotropic population, as described by our model.
@@
In contrast, around the current sheet center, where the magnetic field reverses, the ion VDFs exhibit a perpendicular temperature anisotropy, $T_{i\bot}/T_{i\parallel}>1$, consistent with Speiser-type ion VDFs [nakamura_ion_1998, lottermoser_ion_1998, hietala_ion_2015]. Co-located with this perpendicular temperature anisotropy, the curvature parameter is $\kappa\ll 1$, indicating that ions are following quasi-adiabatic trajectories, such as in Speiser orbits. Using Eq. 9 with the observed $T_{i\bot}/T_{i\parallel}\approx 1.4$, we obtain $n_{s}/n_{i}\approx 0.57$ within the upper range of values reported in previous observations [artemyev_proton_2010]. Hence, the MMS observations indicate that within the current sheet, the ion current is carried by a quasi-adiabatic, perpendicular, anisotropic population, consistent with the Speiser population used in the model.
Figure 4: Distribution $\textrm{PDF}(T_{i\bot}/T_{i\parallel},\beta_{i\parallel})$ in: (a) the near-Earth magnetotail BBFs from MMS [richard_fast_2023], (b) the midtail current sheet [kamaletdinov_thin_2024], (c) PIC-LP simulation of forced magnetic reconnection in a Lembege and Pellat equilibrium magnetotail current sheet [an_configuration_2022], and (d) PIC-HC simulation of forced reconnection in a Harris-like equilibrium current sheet with cold ions [norgren_presence_2021].
@@
To assess whether the theoretical limits bound the observed anisotropy, we present the joint probability density function (PDF) of $\beta_{i\parallel}$ and $T_{i\bot}/T_{i\parallel}$, for the dataset of 235,635 ion VDFs, in Fig. 4a. It was previously suggested that, for this dataset, curvature scattering is responsible for ion isotropization [richard_fast_2023]. We estimate the electron current fraction $\eta_{e}=j_{ey}/j_{0}$, with $j_{ey}$ the electron current calculated using the measured number density and electron bulk velocity and $j_{0}$ calculated using the measured number density and ion temperature, and find a median value of $\eta_{e}=0.3^{+0.4}_{-0.2}$. To estimate the ratio of drift speed of Speiser ions to thermal speed, we compute $\eta_{s}=|\nabla\times\bm{B}|/\mu_{0}j_{0}(n_{s}/n_{i})$, at the center of the current sheet ($B/B_{0}<0.1$), where Speiser ions dominate the current contribution. We compute $|\nabla\times\bm{B}|$ using the curlometer technique [chanteur_spatial_1998, chanteur_spatial_1998a], and adopt a typical fraction of Speiser ions $n_{s}/n_{i}\simeq 0.15$ [artemyev_proton_2010]. It yields, $\eta_{s}=3^{+2}_{-1}$. Estimating the cold ion current fractions, $\eta_{c}$, from the full dataset is challenging due to difficulties in isolating the cold ion population in the ion VDF and in accurately measuring its temperature [li_quantification_2021]. Instead, we adopt representative values of the cold ions perpendicular temperature $T_{c\bot}$ consistent with prior studies, choosing $T_{c\bot}\in[200~\textrm{eV},\,600~\textrm{eV}]$ [li_quantification_2021], and $T_{i\parallel}\in[1~\textrm{keV},\,6~\textrm{keV}]$ measured in the dataset at the current sheet edges $|B|/B_{0}\in[0.2,0.7]$, so that $\eta_{c}\in[1/10,1/2]$. We find that the observed distribution of temperature anisotropy is consistent with the predicted limits for current-sheet stability, particularly in the parallel-anisotropy regime. Although a substantial fraction of the dataset exhibits perpendicular anisotropies that exceed the threshold given by Eq. 12, these cases rarely surpass the limit corresponding to $\kappa=0.1$ (light cyan line), obtained by replacing $\kappa^{2}=1$ with $\kappa^{2}=0.01$ in the derivation of Eq. 12. Even for such values of $\kappa\in[0.1,1]$, ions undergo strong diffusion $\Delta\mathcal{I}_{z}/\mathcal{I}_{z}\sim\kappa$ [buchner_regular_1989]. Hence, in analogy to the thresholds for temperature anisotropy-driven instabilities, the $\kappa=1$ corresponds to the marginal stability, while the $\kappa=0.1$ threshold corresponds to a faster growth rate. Therefore, the MMS observations support well the thresholds derived in Eqs. 8 and 12.

### v2

@@ II.2.2 Speiser dominated regime
In the second regime, we consider that cold ions provide a negligible current ($n_{c}\approx 0$, $j_{cy}\approx 0$), and Speiser ions carry the current. Hence, this regime corresponds to perpendicular temperature anisotropy with
[…]
where we have assumed $T_{i\bot}\approx T_{s\bot}$. Substituting Eq. 8 into Eq. 2, the current density becomes
[…]
where we have used the approximation of the Speiser ion VDF [20, 19, 8], with $A_{s}\approx 1/2$. The corresponding adiabaticity parameter becomes
[…]
To avoid the scattering of hot ions into a quasi-adiabatic Speiser regime (when the entire ion population would move with the drift speed exceeding thermal speed and the current sheet would be strongly unstable), the magnetic field curvature parameter $\kappa$ must remain close to unity [17, 23]. Thus, $\kappa=1$ represents the marginal stability condition for the perpendicular anisotropy. Therefore, from Eq. 10, we obtain the second threshold
[…]
In the limit $\eta_{s}\ll 1$ ($v_{sD}\ll v_{iT\bot}$), the Speiser population is embedded within the hot ion population, and the threshold reduces to $\mathcal{R}_{s}(\eta_{e},\eta_{s}\rightarrow 0,\beta_{i\parallel})\rightarrow 1/\eta_{e}\beta_{i\parallel}$. In this regime, magnetic field stretching is governed by the electron current such that for a given $\beta_{i\parallel}$ and $T_{i\bot}/T_{i\parallel}$, ions with $v_{iT\bot}>V_{Ai}/\sqrt{\eta_{e}}$ will scatter into a Speiser-type orbit. In contrast, for $\eta_{s}\gg 1$ ($v_{sD}\gg v_{iT\bot}$), the threshold approaches $\mathcal{R}_{s}(\eta_{e},\eta_{s}\rightarrow\infty,\beta_{i\parallel})\rightarrow 1$, corresponding to the most unstable configuration with a strong drift current carried by quasi-adiabatic Speiser orbits, characteristic of a drift-type instability [45, 28].
@@ III Spacecraft Observations
We validate our model using in situ spacecraft observations in the Earth’s magnetotail. We employ measurements in the near-Earth magnetotail ($X_{GSM}>-29~R_{E}$, in Geocentric Solar Magnetospheric (GSM) coordinates) from the Magnetospheric Multiscale (MMS) mission [18] and in the mid-tail at lunar distances ($X_{GSM}\sim-60~R_{E}$) from the Acceleration, Reconnection, Turbulence, and Electrodynamics of the Moon’s Interaction with the Sun (ARTEMIS) mission [4].
@@ III.1 Near-Earth reconnection
The MMS dataset includes 516 fast plasma flows measured between 2017 and 2021, at $X_{GSM}\geq-29~R_{E}$, in the central plasma sheet corresponding to $\beta_{i}\geq 0.5$, where $\beta_{i}=2\mu_{0}k_{B}T_{i}/B^{2}$ is the ion plasma beta [62]. Magnetic field measurements are obtained from the fluxgate magnetometer [68], while ion and electron VDFs and their moments are provided by the Fast Plasma Investigation instrument [58]. We correct the ion VDFs for penetrating radiation using the empirical background model described in Ref. 36. To improve counting statistics and reduce uncertainties in moment calculations [37], we average three consecutive ion VDFs over a 450 ms window. It yields a total of 235,635 ion VDFs for analysis.
To illustrate the limiting cases of cold ion-dominated regime, corresponding to parallel anisotropy, and Speiser ion-dominated, corresponding to perpendicular anisotropy, we examine a reconnection outflow observed by MMS on 21 August 2017[65] [Fig. 2]. During this interval, the magnetic field in Fig. 2(a) displays large-amplitude oscillations attributed to a flapping motion of the magnetotail current sheet [63], so that MMS samples across the reconnection outflow shown in Fig. 2(b). Here, we rotated the magnetic field and the ion bulk velocity to the $\bm{\hat{l}}\bm{\hat{m}}\bm{\hat{n}}$ coordinates system, where $\bm{\hat{l}}=\left[0.97,\,-0.25\,0.02\right]\,\textrm{GSM}$ is the maximum variance direction obtained from maximum variance analysis (MVA) of the magnetic field with $\lambda_{max}/\lambda_{int}=15$, $\bm{\hat{n}}=(\bm{\hat{l}}\times\bm{j})/|(\bm{\hat{l}}\times\bm{j})|$, and $\bm{\hat{m}}=\bm{\hat{n}}\times\bm{\hat{l}}$ [69]. In this coordinate system, at $|B_{l}|\approx 0$, MMS is around the center of the current sheet, while maxima in $|B_{l}|$ correspond to the current sheet edges. We present the ion temperature anisotropy $T_{i\bot}/T_{i\parallel}$ [Fig. 2(d)] with respect to the local magnetic field, where we calculated $T_{i\bot}$ and $T_{i\parallel}$ rotating the temperature tensor $\mathbf{T}_{i}$ in the magnetic field-aligned (MFA) frame [71]. In this frame, one axis aligns with the local magnetic field direction $\hat{\bm{b}}$. The perpendicular directions are set to equalize the remaining diagonal components of the ion temperature tensor $\mathbf{T}_{i}=T_{i\parallel}\hat{\bm{b}}\hat{\bm{b}}+T_{i\bot}(\mathbf{1}-\hat{\bm{b}}\hat{\bm{b}})+\bm{\Theta}_{i}$, where $T_{i\parallel}$ and $T_{i\bot}$ are the parallel and perpendicular temperatures, and $\bm{\Theta}_{i}$ is the agyrotropic part with zero diagonal. We compute the magnetic field curvature radius $R_{c}$, and the corresponding adiabaticity parameter $\kappa$, using the multi-spacecraft curlometer technique [21, 22]. We find that the temperature anisotropy [Fig. 2(d)] and the adiabaticity parameter [Fig. 2(e)] vary strongly across the reconnection outflow and are inversely correlated, indicating interplay between magnetic field curvature and the ion VDFs.
@@
To examine the spatial variation of the temperature anisotropy across the current sheet, we construct profiles of $T_{i\parallel}$ and $T_{i\bot}$ as functions of the reconnecting magnetic field component $B_{l}$ [Figs. 3(a)- 3(d)]. We normalize $B_{l}$ by the background (lobe) magnetic field strength $B_{0}=B\sqrt{1+\beta_{i}}$, estimated from pressure balance of the magnetotail plasmasheet [13]. To facilitate comparison with numerical simulation (see Section IV), we normalize the ion temperatures $T_{i}$, $T_{i\parallel}$, $T_{i\bot}$ to the electromagnetic energy available per particle $m_{i}V_{A0}^{2}$, where $V_{A0}=B_{0}/\sqrt{\mu_{0}m_{i}n_{0}}$ is the Alfvén speed and $n_{0}$ the ion density within the current sheet. For the flapping event presented in Fig. 2, we obtain $B_{0}=21.6\,\mathrm{nT}$ and $n_{0}=0.27\,\mathrm{cm}^{-3}$, the latter computed as the average number density over intervals with $|B_{l}|/B_{0}<0.1$ well within the current sheet.
@@
Figure 3: Comparison between MMS observations and PIC simulation results for ion temperature anisotropy. Panels (a)–(d) and (e)–(h) show 1D profiles of: (a, e) ion temperature parallel to the local magnetic field, (b, f) ion temperature perpendicular to the magnetic field, (c, g) ion temperature anisotropy, and (d, h) the magnetic field curvature parameter. Black crosses represent MMS measurements during the event shown in Fig. 2. Solid blue and red lines correspond to profiles extracted along magnetic field lines in the PIC simulation, intersecting the $z=0$ plane at $\delta x=5d_{i}$ and $\delta x=11d_{i}$, respectively. Panels (i) and (j) present 2D maps of ion temperature anisotropy and curvature parameter from the PIC simulation. The black contour indicates $\kappa=1$. The black, thick-dashed lines indicate the separatrices. The green contour indicates the region of interest for Fig. 4(c).
@@
At the edges of the current sheet – corresponding to where $B_{l}/B_{0}=\pm 0.3$ – the ion VDFs are elongated parallel to the magnetic field, with $T_{i\bot}/T_{i\parallel}<1$, characteristic of cold, field-aligned counter streaming beams [41, 40] as described in our model. We find that instances of parallel anisotropy coincide with $\kappa>1$, indicating that the cold ions follow regular adiabatic orbits. These observations suggest that outside the current sheet center, the ion VDF primarily consists of an adiabatic, parallel-anisotropic population, as described by our model.
@@
In contrast, around the current sheet center, where the magnetic field reverses, the ion VDFs exhibit a perpendicular temperature anisotropy, $T_{i\bot}/T_{i\parallel}>1$, consistent with Speiser-type ion VDFs [53, 49, 40]. Co-located with this perpendicular temperature anisotropy, the curvature parameter is $\kappa\ll 1$, indicating that ions are following quasi-adiabatic trajectories, such as in Speiser orbits. Using Eq. 8 with the observed $T_{i\bot}/T_{i\parallel}\approx 1.4$, we obtain $n_{s}/n_{i}\approx 0.57$ within the upper range of values reported in previous observations [10]. Hence, the MMS observations indicate that within the current sheet, the ion current is carried by a quasi-adiabatic, perpendicular, anisotropic population, consistent with the Speiser population used in the model.
@@
To assess whether the theoretical limits (Eqs. 7 and 11) bound the observed anisotropy, we present the joint probability density function (PDF) of $\beta_{i\parallel}$ and $T_{i\bot}/T_{i\parallel}$, for the dataset of 235,635 ion VDFs, in Fig. 4a. It was previously suggested that, for this dataset, curvature scattering is responsible for ion isotropization [65]. We estimate the electron current fraction $\eta_{e}=j_{ey}/j_{0}$, with $j_{ey}$ the electron current calculated using the measured number density and electron bulk velocity, and $j_{0}$ calculated using the measured number density and ion temperature. We find a median value of $\eta_{e}=0.3$ with a $95\%$ confidence interval $\eta_{e}\in[0.1,0.7]$. To estimate the ratio of drift speed of Speiser ions to thermal speed, we compute $\eta_{s}=|\nabla\times\bm{B}|/\mu_{0}j_{0}(n_{s}/n_{i})$, at the center of the current sheet ($B/B_{0}<0.1$), where Speiser ions dominate the current contribution. We compute $|\nabla\times\bm{B}|$ using the curlometer technique [21, 22], and adopt a typical fraction of Speiser ions $n_{s}/n_{i}\simeq 0.15$ [10]. It yields a median value of $\eta_{s}=3.4$ with a $95\%$ confidence interval of $\eta_{s}\in[2.3,4.8]$. Estimating the cold ion current fractions, $\eta_{c}$, from the full dataset is challenging due to difficulties in isolating the cold ion population in the ion VDF and in accurately measuring its temperature [47]. Instead, we adopt representative values of the cold ions perpendicular temperature $T_{c\bot}$ consistent with prior studies, choosing $T_{c\bot}\in[200~\textrm{eV},\,600~\textrm{eV}]$ [47], and $T_{i\parallel}\in[1~\textrm{keV},\,6~\textrm{keV}]$ measured in the dataset at the current sheet edges $|B|/B_{0}\in[0.2,0.7]$, so that $\eta_{c}\in[0.1,0.5]$. We find that the observed distribution of temperature anisotropy is consistent with the predicted limits for current-sheet stability, particularly in the parallel-anisotropy regime. Although a substantial fraction of the dataset exhibits perpendicular anisotropies that exceed the threshold given by Eq. 11, these cases rarely surpass the limit corresponding to $\kappa=0.1$ (light cyan line), obtained by replacing $\kappa^{2}=1$ with $\kappa^{2}=0.01$ in the derivation of Eq. 11. Even for such values of $\kappa\in[0.1,1]$, ions undergo strong diffusion $\Delta\mathcal{I}_{z}/\mathcal{I}_{z}\sim\kappa$ [17]. Hence, in analogy to the thresholds for temperature anisotropy-driven instabilities, the $\kappa=1$ corresponds to the marginal stability, while the $\kappa=0.1$ threshold corresponds to a faster growth rate. Therefore, the MMS observations support well the thresholds derived in Eqs. 7 and 11.
Figure 4: Distribution $\textrm{PDF}(T_{i\bot}/T_{i\parallel},\beta_{i\parallel})$ in: (a) the near-Earth magnetotail BBFs from MMS [65], (b) the midtail current sheet [42], (c) PIC-LP simulation of forced magnetic reconnection in a Lembege and Pellat equilibrium magnetotail current sheet [3], and (d) PIC-HC simulation of forced reconnection in a Harris-like equilibrium current sheet with cold ions [55]. The dark blue and yellow dashed and dash–dotted lines show the thresholds given by Eq. 11 for the extreme values of the confidence intervals of $\eta_{s}$ and $\eta_{e}$, respectively. The red dashed and dash–dotted lines indicate the thresholds from Eq. 7 for the extreme values of the confidence interval of $\eta_{c}$. Solid cyan and pink lines mark the thresholds corresponding to the nominal values $(\eta_{e}=0.4,~\eta_{s}=3)$, and $\eta_{c}=0.2$, respectively. The translucent cyan line denotes the threshold for perpendicular temperature anisotropy modified for $\kappa=0.1$.